# Normal forms for ordinary differential operators, I

### v1

@@ 2.3 Homogeneous canonical polynomials
  3. 3.
    Now we have: $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a HCP. Moreover, if $Sdeg_{A}(M)\neq-\infty$, then $Sdeg_{A}(H_{0})=Sdeg_{A}(M)+1$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. And if $Sdeg_{A}(M)=-\infty$, then $Sdeg_{A}(H_{0})=-\infty$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. Since our original solution $H\in\hat{D}_{1}^{sym}$, we have $\mathop{\rm\bf ord}(H)<\infty$ and therefore $\mathop{\rm\bf ord}(c_{i})<\infty$ for all $i$, i.e. $c_{i}\in\tilde{K}[\tilde{\partial}^{-1},\tilde{\partial}]$ and $H$ is a HCPC.
@@
  1. 1.
    $H=b_{i}B_{i}D^{u},M=c_{j}B_{j}D^{v}$, where $i\geq 1-u$ if $u<0$ and $j\geq 1-v$ if $v<0$, $b_{i},c_{j}\in\tilde{K}$. We have by lemma 2.5
@@
  2. 2.
    $H=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $M=a_{l,m}\Gamma_{m}A_{j}D^{v}$, $b_{i},a_{l,m}\in\tilde{K}$. By Lemma 2.5 item 12 we know $B_{i}D^{u}=D^{u}B_{i+u}$, hence $i\geq 1-u$, otherwise $H=0$. Then we have
@@
    $B_{i}D^{u}\Gamma_{m}A_{j}D^{v}=B_{i}(x\partial+u)^{m}D^{u}A_{j}D^{v}=\xi^{uj}B_{i}(x\partial+u)^{m}A_{j}D^{u}D^{v}\\ =\xi^{uj}B_{i}A_{j}(x\partial+u)^{m}D^{u}D^{v}=\xi^{(u+i-1)j}B_{i}(x\partial+u)^{m}D^{u}D^{v}\\ =\xi^{(u+i-1)j}(i-1+u)^{m}B_{i}D^{u}D^{v}$
@@
  3. 3.
    $H=a_{l,m}\Gamma_{m}A_{j}D^{v},M=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $b_{i},a_{l,m}\in\tilde{K}$. We have
@@
  4. 4.
    $H=a_{i,m}\Gamma_{m}A_{i}D^{u},M=a_{j,n}\Gamma_{n}A_{j}D^{v}$, $a_{i,m},a_{j,n}\in\tilde{K}$. We have
@@
    $\Gamma A_{i}D^{u}\Gamma_{n}A_{j}D^{v}=\Gamma_{m}A_{i}(x\partial+u)^{u}D^{u}A_{j}D^{v}=\xi^{uj}A_{i}\Gamma_{m}(x\partial+u)^{n}A_{j}D^{u}D^{v}\\ =\xi^{uj}A_{i+j}\Gamma_{m}(x\partial+u)^{n}D^{u}D^{v}$
@@
  1. 1.
    Consider $Sdeg_{A}(T)$. Observe that in cases 1-4 our claim is true. So, it is true for a product of any two HCPCs. Note also that in the case 4, for $H,M$ being monomials, we have the equality $Sdeg_{A}(T)=Sdeg_{A}(H)+Sdeg_{A}(M)$, however, in general, if we take the product of sums of monomials, a strict inequality can appear.
  2. 2.
    Consider $Sdeg_{B}(T)$. Then explicit formulae of cases 1-4 show our assertion in general case.
@@
  2. 2.
    Suppose $\tilde{H}$ is the result we get in step 1. From the discussion in step 1, we know
@@ 2.4 Some necessary conditions on the Schur operator
$H\partial^{u}=\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}(1-B_{1}-\cdots-B_{u})\\ =\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}-\sum_{j=1}^{u}\sum_{i=0}^{k-1}\sum_{m=0}^{d}\xi^{i(j-1)}(j-1)^{m}a_{m,i}B_{j}$
@@
$(\alpha_{(d+1)k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{(d+1)k-k+1})(u-1)^{d}+(\cdots)(u-1)^{d-1}\\ +\cdots+(\alpha_{k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{1})=0$
@@
$(\alpha_{(d+1)k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{(d+1)k-k+1})(u-2)^{d}+(\cdots)(u-2)^{d-1}\\ +\cdots+(\alpha_{k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{1})=0$
$(\alpha_{(d+1)k}\xi^{k-1}+\cdots+\alpha_{(d+1)k-k+1})+(\cdots)\\ +\cdots+(\alpha_{k}\xi^{k-1}+\cdots+\alpha_{1})=0$
@@ 2.5 Some necessary conditions on normal forms
Let $Q\in D_{1}$ be a normalized operator as in the previous section. Let $P\in D_{1}$ be a monic operator of positive order.
@@
For a given pair of monic operators $Q,P\in D_{1}$ with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q\geq 0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p\geq 0$ we define a normal form of $P$ with respect to $Q$ as the operator $P^{\prime}:=S^{-1}PS$, where $S$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$.
@@
From previous section we know that $S,S^{-1}$ satisfy condition $A_{q}(0)$, where $S$ is any monic Schur operator for normalised $Q\in D_{1}$.
@@
$P_{-1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}\Gamma_{m}D^{-1}\\ =\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m}D^{-1}\in D_{1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m-1}x\in D_{1}$
@@
Suppose $Q\in D_{1}$ is a monic operator with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$. Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{q}(0)$, $\mathop{\rm\bf ord}(P)=p$. Put $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5.
@@ 3 Normal forms for commuting operators
In this section we give a convenient description of the centraliser $C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ and of normal forms of the operator $P$ with respect to $Q$. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve (according to the well known classification theory of commuting ordinary differential operators such sheaves describe rank one subrings of commuting operators with a given spectral curve, cf. [43, Th. 10.26]).
@@
Now consider the skew polynomial ring ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$, where $\sigma(a_{0},\ldots,a_{k-1})=(a_{k-1},a_{0},\ldots,a_{k-2})$  [^5] We use a standard notations and constructions from the books [9] and [23]. A short self contained exposition of all necessary constructions and facts see e.g. in [43], Ch. 2,3 and 13.1. Recall that this notation means that we have the following commutation relation between $D$ and $(a_{0},\ldots,a_{k-1})$: $(a_{0},\ldots,a_{k-1})D=D\sigma(a_{0},\ldots,a_{k-1})$. . It is a right and left noetherian ring. The multiplicatively closed subset $S=\{D^{k},k\geq 0\}$ obviously satisfies the right Ore condition, consists of regular elements, and ${\bf ass}(S)=0$. So, the right quotient ring $\mathfrak{B}={{\cal R}}_{S}$ exists. Clearly,
@@
Let $C(\mathfrak{B})$ be the center of $\mathfrak{B}$, Obviously, $D^{k},D^{-k}\in C(\mathfrak{B})$. Since an element $(h_{0},h_{1},\dots,h_{k-1})\in\tilde{K}^{\oplus k}$ doesn’t commute with any $D^{l}$, if not all of $h_{i}$ equal to each other, we have
@@
Therefore, for any $P\in C(\partial^{k})$ the characteristic polynomial $\det(\psi(P)-\lambda)\in\tilde{K}[\lambda,D^{k}]$ defines an algebraic relation between $P$ and $\partial^{k}$ or, in degenerate cases, it defines an algebraic dependence over $\tilde{K}$. Note that if $P^{\prime}$ is a normal form of a monic differential operator $P$ with respect to the commuting with $P$ differential operator $Q$, and the order of $P$ is coprime with the order of $Q$, then $P^{\prime}$ is monic and $\det(\psi(P)-\lambda)\in\tilde{K}[\lambda,D^{k}]$ is a polynomial of the form similar to the Burchnall-Chaundy polynomial [^7] Recall the famous Burchnall-Chaundy lemma: if $P,Q\in D_{1}$ are differential operators of coprime orders $p$ and $q$, then they are algebraically dependent, and satisfy a polynomial relation of the form $f(X,Y)=\alpha X^{q}\pm Y^{p}+\ldots=0$, where $GCD(p,q)=1$ and $\lambda\neq 0$ (here the weighted degree of $f$ is $pq$, and $\ldots$ mean terms of lower weighted degree, where the weight of $X$ is $p$, and the weight of $Y$ is $q$; in particular, for coprime $p$ and $q$ it is automatically irreducible). , i.e. $\det(\psi(P)-\lambda)=\pm\lambda^{q}\pm D^{kp}+\ldots$, in particular it is irreducible and therefore coincides with the Burchnall-Chaundy polynomial for $P$ and $Q$ up to a multiplicative constant.
Embedding normal forms to the ring $E_{q}$, we can transform them to a simpler form by conjugation:
@@
Assume $\tilde{P}\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{q}$ is a monic operator with $\mathop{\rm ord}\nolimits_{\tilde{D}}(\tilde{P})=p$. Let $p=dn$, $q=dm$, where $d=GCD(p,q)$.
Then there exists a monic invertible operator $S\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{q}$ with $\mathop{\rm ord}\nolimits_{\tilde{D}}(S)=0$ such that all coefficients of the operator $S^{-1}\tilde{P}S$ commute with $\tilde{D}^{d}$.
@@
We will find the operator $S$ as the limit of a Cauchy sequence888with respect to the topology defined by the pseudo-valuation $-\mathop{\rm ord}\nolimits_{\tilde{D}}$. Assume
@@
Then there is a normalised normal form $P^{\prime}$ of $P$ with respect to $Q$ uniquely defined up to conjugation by an element from the centraliser $C(\partial^{q})$. Namely, we can explicitly describe its image under the embedding $\hat{\Phi}$ as follows:
@@
Let’s call coefficients $p_{l,j}$ supplementary to the list above as coordinates of $P^{\prime}$. If $P^{\prime}$ and $P^{\prime\prime}$ are two operators with different values of coordinates, then there are no invertible operators $S\in C(\partial^{q})$ such that $P^{\prime}=S^{-1}P^{\prime\prime}S$.
@@
On the other hand, since $\tilde{P}^{\prime\prime}\in C(\partial^{q})$, we know from lemma 2.4 that always $\tilde{p}_{l,j}=$ for $j=0,\ldots,-l-1$ if $l<0$ (in particular, the number of zero coefficients for $p-q<l<0$ in case $p<q$ is constant). Taking $S=\prod_{j=1}^{q-1}S_{j}$ we get the needed operator: the normal form $S^{-1}P^{\prime}S$ will satisfy all conditions as stated. This completes the proof of the first statement.
@@
([43, Th. 10.26]) There is a one-to-one correspondence
$[B\subset{D}\mbox{\quad of rank $1$}]/\thicksim\longleftrightarrow[(C,p,{{\cal F}})\mbox{\quad of rank $1$}]/\simeq$
@@
  • $[B]$ means a class of equivalent commutative elliptic subrings (i.e. $B$ containing a monic differential operator), where $B\sim B^{\prime}$ iff $B=f^{-1}B^{\prime}f$, $f\in D^{*}$.
@@
The self contained proof of this theorem in such a form is given in [43, Ch. 9, 10], and it uses two other correspondences between these data and equivalence classes of Schur pairs (we recall the definition below, as they will play a key role in our statements). The proof given in [43] is an elaborated version of Mulase’s proof from [26] in a spirit of works [29], [28] and their higher dimensional generalisations in [40], [17].
@@
Let $B\subset D_{1}$ be an elliptic subring. Then by Schur theory from [33], cf. [43, T. 4.6, C. 4.7], there exists an invertible operator $S=s_{0}+s_{1}\partial^{-1}+\ldots$ in the usual (Schur’s) ring of pseudo-differential operators $E=K[[x]]((\partial^{-1}))$ such that $A:=S^{-1}BS\subset K((\partial^{-1}))$. Consider the homomorphism (of vector spaces)
@@
$1\circ:\tilde{K}[A_{1}]((\tilde{D}^{-1}))\subset E_{q}\rightarrow\tilde{K}((\tilde{D}^{-1})),\quad\sum_{l}p_{l}\tilde{D}^{l}\mapsto\sum_{l}p_{l,0}\tilde{D}^{l}$
(cf. lemma 3.3) defines a structure of a $\tilde{K}[A_{1}]((\tilde{D}^{-1}))$-module on the space $\tilde{K}((\tilde{D}^{-1}))$: for any $P\in\tilde{K}((\tilde{D}^{-1}))$ and $Q\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ we put $P\cdot Q=1\circ(PQ)$.
@@
Analogously, we can define the space $W^{\prime}:=F^{\prime}\cdot S\subset\tilde{K}((\tilde{D}^{-1}))$, where $F^{\prime}=\tilde{K}[\tilde{D}]$ and $S\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ is an operator from lemma 3.3. If $B^{\prime}=\hat{\Phi}(S^{-1}BS)$, where $S$ is Schur operator from proposition 2.2, and $A^{\prime}:=S^{-1}B^{\prime}S$, where $S$ is an operator from corollary 3.1, then, clearly, the modules $F^{\prime}$ and $F$ are isomorphic ($F^{\prime}$ as a $B^{\prime}$-module, $F$ as a $B$-module), and $W^{\prime}$ and $F$ are isomorphic ($W^{\prime}$ is a $B^{\prime}$-module, $F$ as a $B$-module), so also $W^{\prime}\simeq W$. Moreover, all these modules are isomorphic as filtered modules (with respect to the order filtration).
For subrings in $K((z))$ we can introduce the same notion of rank as for subrings in $D_{1}$999here $z$ will play a role of $\partial^{-1}$ or $\tilde{D}^{-1}$:

### v3

@@ 2. We have two possible case of such monomials:
3. Now we have: $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a HCP. Moreover, if $Sdeg_{A}(M)\neq-\infty$, then $Sdeg_{A}(H_{0})=Sdeg_{A}(M)+1$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. And if $Sdeg_{A}(M)=-\infty$, then $Sdeg_{A}(H_{0})=-\infty$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. Since our original solution $H\in\hat{D}_{1}^{sym}$, we have $\mathop{\rm\bf ord}(H)<\infty$ and therefore $\mathop{\rm\bf ord}(c_{i})<\infty$ for all $i$, i.e. $c_{i}\in\tilde{K}[\tilde{\partial}^{-1},\tilde{\partial}]$ and $H$ is a HCPC.
@@
1. $H=b_{i}B_{i}D^{u},M=c_{j}B_{j}D^{v}$, where $i\geq 1-u$ if $u<0$ and $j\geq 1-v$ if $v<0$, $b_{i},c_{j}\in\tilde{K}$. We have by lemma 2.5
@@
2. $H=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $M=a_{l,m}\Gamma_{m}A_{j}D^{v}$, $b_{i},a_{l,m}\in\tilde{K}$. By Lemma 2.5 item 12 we know $B_{i}D^{u}=D^{u}B_{i+u}$, hence $i\geq 1-u$, otherwise $H=0$. Then we have
@@
$B_{i}D^{u}\Gamma_{m}A_{j}D^{v}=B_{i}(x\partial+u)^{m}D^{u}A_{j}D^{v}=\xi^{uj}B_{i}(x\partial+u)^{m}A_{j}D^{u}D^{v}\\ =\xi^{uj}B_{i}A_{j}(x\partial+u)^{m}D^{u}D^{v}=\xi^{(u+i-1)j}B_{i}(x\partial+u)^{m}D^{u}D^{v}=\xi^{(u+i-1)j}(i-1+u)^{m}B_{i}D^{u}D^{v}$
@@
3. $H=a_{l,m}\Gamma_{m}A_{j}D^{v},M=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $b_{i},a_{l,m}\in\tilde{K}$. We have
@@
4. $H=a_{i,m}\Gamma_{m}A_{i}D^{u},M=a_{j,n}\Gamma_{n}A_{j}D^{v}$, $a_{i,m},a_{j,n}\in\tilde{K}$. We have
@@
$\Gamma A_{i}D^{u}\Gamma_{n}A_{j}D^{v}=\Gamma_{m}A_{i}(x\partial+u)^{u}D^{u}A_{j}D^{v}=\xi^{uj}A_{i}\Gamma_{m}(x\partial+u)^{n}A_{j}D^{u}D^{v}=\xi^{uj}A_{i+j}\Gamma_{m}(x\partial+u)^{n}D^{u}D^{v}$
@@
1. Consider $Sdeg_{A}(T)$. Observe that in cases 1-4 our claim is true. So, it is true for a product of any two HCPCs. Note also that in the case 4, for $H,M$ being monomials, we have the equality $Sdeg_{A}(T)=Sdeg_{A}(H)+Sdeg_{A}(M)$, however, in general, if we take the product of sums of monomials, a strict inequality can appear.
2. Consider $Sdeg_{B}(T)$. Then explicit formulae of cases 1-4 show our assertion in general case.
@@ 1. Assume
2. Suppose $\tilde{H}$ is the result we get in step 1. From the discussion in step 1, we know
@@ 2.4 Some necessary conditions on the Schur operator
$H\partial^{u}=\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}(1-B_{1}-\cdots-B_{u})=\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}-\sum_{j=1}^{u}\sum_{i=0}^{k-1}\sum_{m=0}^{d}\xi^{i(j-1)}(j-1)^{m}a_{m,i}B_{j}$
@@
$(\alpha_{(d+1)k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{(d+1)k-k+1})(u-1)^{d}+(\cdots)(u-1)^{d-1}+\cdots+(\alpha_{k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{1})=0$
@@
$(\alpha_{(d+1)k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{(d+1)k-k+1})(u-2)^{d}+(\cdots)(u-2)^{d-1}+\cdots+(\alpha_{k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{1})=0$
$(\alpha_{(d+1)k}\xi^{k-1}+\cdots+\alpha_{(d+1)k-k+1})+(\cdots)+\cdots+(\alpha_{k}\xi^{k-1}+\cdots+\alpha_{1})=0$
@@ 2.5 Some necessary conditions on normal forms
Let $Q\in D_{1}$ be a normalized operator as in the previous section.
@@
For a given pair of monic operators $Q,P\in D_{1}$ with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p\geq 0$ we define a normal form of $P$ with respect to $Q$ as the operator $P^{\prime}:=S^{-1}PS$, where $S$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$.
@@
From previous section we know that $S,S^{-1}$ satisfy condition $A_{q}(0)$, where $S$ is any monic Schur operator for normalised $Q\in D_{1}$. Moreover, if $S$ is a monic operator from $\hat{D}_{1}^{sym}$ of order zero satisfying condition $A_{q}(0)$ (but not necessary with $S_{-1}=0$), then $S^{-1}$ also satisfies the same properties. The proof is just the same as in corollary 2.3.
@@
$P_{-1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}\Gamma_{m}D^{-1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m}D^{-1}\in D_{1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m-1}x\in D_{1}$
@@
Suppose $Q\in D_{1}$ is a normalized operator with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$. Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{q}(0)$, $\mathop{\rm\bf ord}(P)=p$. Put $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5.
@@ 3 Normal forms for commuting operators
In this section we give a convenient description of the centraliser $C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ and of normal forms of the operator $P$ with respect to $Q$. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve (according to the well known classification theory of commuting ordinary differential operators such sheaves describe rank one subrings of commuting operators with a given spectral curve, cf. [47, Th. 10.26]).
@@
Now consider the skew polynomial ring ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$, where
$\sigma(a_{0},\ldots,a_{k-1})=(a_{k-1},a_{0},\ldots,a_{k-2})$  [^4] We use a standard notations and constructions from the books [13] and [27]. A short self contained exposition of all necessary constructions and facts see e.g. in [47], Ch. 2,3 and 13.1. Recall that this notation means that we have the following commutation relation between $D$ and $(a_{0},\ldots,a_{k-1})$: $(a_{0},\ldots,a_{k-1})D=D\sigma(a_{0},\ldots,a_{k-1})$. . It is a right and left noetherian ring. The multiplicatively closed subset $S=\{D^{k},k\geq 0\}$ obviously satisfies the right Ore condition, consists of regular elements, and ${\bf ass}(S)=0$. So, the right quotient ring $\mathfrak{B}={{\cal R}}_{S}$ exists. Clearly,
@@
Let $C(\mathfrak{B})$ be the center of $\mathfrak{B}$. Obviously, $D^{k},D^{-k}\in C(\mathfrak{B})$. Since an element $(h_{0},\ldots,h_{k-1})\in\tilde{K}^{\oplus k}$ doesn’t commute with any $D^{l}$, if not all of $h_{i}$ equal to each other, we have
@@
Therefore, for any $P\in C(\partial^{k})$ the characteristic polynomial $\det(\psi(\hat{\Phi}(P))-\lambda)\in\tilde{K}[\lambda,D^{k}]$ defines an algebraic relation between $P$ and $\partial^{k}$ or, in degenerate cases, it defines an algebraic dependence over $\tilde{K}$. Note that if $P^{\prime}$ is a normal form of a monic differential operator $P$ with respect to the commuting with $P$ differential operator $Q$, and the order of $P$ is coprime with the order of $Q$, then $P^{\prime}$ is monic and $\det(\psi(\hat{\Phi}(P))-\lambda)\in\tilde{K}[\lambda,D^{k}]$ is a polynomial of the form similar to the Burchnall-Chaundy polynomial, i.e. $\det(\psi(P)-\lambda)=\pm\lambda^{q}\pm D^{kp}+\ldots$, in particular it is irreducible and therefore coincides with the Burchnall-Chaundy polynomial for $P$ and $Q$ up to a multiplicative constant.
Remark 3.4.
Elements of the centralizer $C(\partial^{k})$ embedded to $\mathfrak{B}$ we’ll call as a vector form presentation, and elements embedded to $M_{k}(\tilde{K}[D^{k}])$ – as a matrix form presentation.
Translating the description of the centralizer $C(\partial^{k})$ given in proposition 2.1, item 2, into the vector form, we get that $\hat{\Phi}(C(\partial^{k}))$ consists of Laurent polynomials in $\tilde{D}$ with coefficients from $K^{\oplus k}$ and with additional conditions: the coefficient $s_{i}$ at $\tilde{D}^{-i}$, $i>0$, has a shape $s_{i,j}=0$ for $j=0,\ldots,i-1$.
Embedding normal forms to the ring $E_{k}$, we can transform them to a simpler form by conjugation:
@@
Assume $\tilde{P}\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{k}$ is a monic operator with $\mathop{\rm ord}\nolimits_{\tilde{D}}(\tilde{P})=p$. Let $p=dn$, $k=q=dm$, where $d=GCD(p,q)$.
Then there exists a monic invertible operator $S\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{k}$ with $\mathop{\rm ord}\nolimits_{\tilde{D}}(S)=0$ such that all coefficients of the operator $S^{-1}\tilde{P}S$ commute with $\tilde{D}^{d}$.
@@
We will find the operator $S$ as the limit of a Cauchy sequence666with respect to the topology defined by the pseudo-valuation $-\mathop{\rm ord}\nolimits_{\tilde{D}}$. Assume
@@
Remark 3.5.
Clearly, if $\tilde{P}\in K^{\oplus k}((\tilde{D}^{-1}))$, then the operator $S$ from lemma 3.3 also belongs to $K^{\oplus k}((\tilde{D}^{-1}))$, cf. remark 3.2.
@@
Then there is a normalised normal form $P^{\prime}$ of $P$ with respect to $Q$ uniquely defined in the conjugacy class by elements from the centraliser $C(\partial^{q})$. Namely, we can explicitly describe its image under the embedding $\hat{\Phi}$ as follows:
@@
Let’s call coefficients $p_{l,j}$ supplementary to the list above as coordinates of $P^{\prime}$. If $P^{\prime}$ and $P^{\prime\prime}$ are two operators with different values of coordinates, then there are no invertible operators $S\in C(\partial^{q})$ such that $P^{\prime}=S^{-1}P^{\prime\prime}S$, and there are no invertible $1\neq S\in C(\partial^{p})$ such that $P^{\prime}=S^{-1}P^{\prime}S$.
@@
On the other hand, since $\tilde{P}^{\prime\prime}\in C(\partial^{q})$, we know from lemma 2.4 that always $\tilde{p}_{l,j}=0$ for $j=0,\ldots,-l-1$ if $l<0$ (in particular, the number of zero coefficients for $p-q<l<0$ in case $p<q$ is constant). Taking $S=\prod_{j=1}^{q-1}S_{j}$ we get the needed operator: the normal form $S^{-1}P^{\prime}S$ will satisfy all conditions as stated. This completes the proof of the first statement.
@@
([47, Th. 10.26]) There is a one-to-one correspondence
$[B\subset{D_{1}}\mbox{\quad of rank $1$}]/\thicksim\longleftrightarrow[(C,p,{{\cal F}})\mbox{\quad of rank $1$}]/\simeq$
@@
  • $[B]$ means a class of equivalent commutative elliptic subrings (i.e. $B$ containing a monic differential operator), where $B\sim B^{\prime}$ iff $B=f^{-1}B^{\prime}f$, $f\in D_{1}^{*}=K[[x]]^{*}$.
@@
The self contained proof of this theorem in such a form is given in [47, Ch. 9, 10], and it uses two other correspondences between these data and equivalence classes of Schur pairs (we recall the definition below, as they will play a key role in our statements). The proof given in [47] is an elaborated version of Mulase’s proof from [30] in a spirit of works [33], [32] and their higher dimensional generalisations in [44], [21].
@@
Let $B\subset D_{1}$ be an elliptic subring. Then by Schur theory from [37], cf. [47, T. 4.6, C. 4.7], there exists an invertible operator $S=s_{0}+s_{1}\partial^{-1}+\ldots$ in the usual (Schur’s) ring of pseudo-differential operators $E=K[[x]]((\partial^{-1}))$ such that $A:=S^{-1}BS\subset K((\partial^{-1}))$. Consider the homomorphism (of vector spaces)
@@
$1\circ:\tilde{K}[A_{1}]((\tilde{D}^{-1}))\simeq\tilde{K}^{\oplus q}((\tilde{D}^{-1}))\subset E_{q}\rightarrow\tilde{K}((\tilde{D}^{-1})),\quad\sum_{l}p_{l}\tilde{D}^{l}\mapsto\sum_{l}p_{l,0}\tilde{D}^{l}$
(cf. lemma 3.3) defines a structure of a $\tilde{K}[A_{1}]((\tilde{D}^{-1}))$-module on the space $\tilde{K}((\tilde{D}^{-1}))$: for any $P\in\tilde{K}((\tilde{D}^{-1}))$ and $Q\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ we put $P\cdot Q=1\circ(PQ)$. Analogously, $K((\tilde{D}^{-1}))\subset\tilde{K}((\tilde{D}^{-1}))$ is a right $K^{\oplus q}((\tilde{D}^{-1}))$-module.
@@
Analogously, we can define the space $W^{\prime}:=F^{\prime}\cdot S\subset\tilde{K}((\tilde{D}^{-1}))$, where $F^{\prime}=\tilde{K}[\tilde{D}]$ and $S\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ is an operator from lemma 3.3 or $W^{\prime\prime}:=F^{\prime\prime}\cdot S\subset K((\tilde{D}^{-1}))$, where $F^{\prime\prime}=K[\tilde{D}]$ if $S\in K^{\oplus q}((\tilde{D}^{-1}))$.
If $B^{\prime}=\hat{\Phi}(S^{-1}BS)$, where $S$ is Schur operator from proposition 2.2 (constructed for some monic differential operator from $B$), and $A^{\prime}:=S^{-1}B^{\prime}S$, where $S$ is an operator from lemma 3.3 (constructed for some operator operator $P^{\prime}\in B^{\prime}$), and, by remark 3.5 $S\in K^{\oplus q}((\tilde{D}^{-1}))$, so that $A^{\prime}\subset K((\tilde{D}^{-1}))$, then, clearly, the modules $F^{\prime\prime}$ and $F$ are isomorphic ($F^{\prime\prime}$ as a $B^{\prime}$-module, $F$ as a $B$-module), and $W^{\prime\prime}$ and $F$ are isomorphic ($W^{\prime\prime}$ is a $A^{\prime}$-module, $F$ as a $B$-module), so also $W^{\prime\prime}\simeq W$. Moreover, all these modules are isomorphic as filtered modules (with respect to the order filtration).
For subrings in $K((z))$ we can introduce the same notion of rank as for subrings in $D_{1}$777here $z$ will play a role of $\partial^{-1}$ or $\tilde{D}^{-1}$: